Suppose consumer spending is specified as C = k50,000 + 0.50Y , investment spending k75,000 and next exports k15,000.

a) Plot a line representing consumer spending and label it C. Also plot a line representing consumer, investment and net exports spending and label it (C+I+Xn). Draw a 45
1
Expert's answer
2021-09-01T12:06:39-0400

The equilibrium level of output is at point E where


AD=AS


At full employment level.



At equilibrium level


AD=AS

By putting values given the question


"Y=AD\\



Y=C+I+X_n\\



Y=k50000+0.50Y+k75000+k15000\\



Y-0.50Y=k140000\\



Y=k\frac{140000}{0.50}\\



Y=k280000"


It is the equilibrium level.